OpenStudy (anonymous):

Find the sum of the roots of 2x^2 = 13x - 15.

OpenStudy (anonymous):

2x^2-13x+15=0 2x^2-15x+2x+15=0 -x(2x-15)+(2x+15)=0 (2x+15)(1-x)=0 2x+15=0 or 1-x=0 x=-15/2 or x=1

OpenStudy (anonymous):

Im not sure that I answer your question

OpenStudy (anonymous):

oh yes, thanks, I got that but then when I checked the answer, it was wrong.

OpenStudy (anonymous):

so is my answer corrcet?

OpenStudy (anonymous):

well, i got the same answer as you did but when i tried to CHECK the answer, it wasn't right.

OpenStudy (anonymous):

doesn't equal out.

OpenStudy (anonymous):

oh I saw the problem

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

2x^2-10x-3x+15=0 2x(x-5)-3(x-5)=0 (2x-3)(x-5)=0 x=3/2 or x=5

OpenStudy (anonymous):

why did you separate them like that though?

OpenStudy (anonymous):

the corrct answer after put #in to check x=5

OpenStudy (anonymous):

just easy for me to see

OpenStudy (anonymous):

okay, thank you :)

OpenStudy (anonymous):

you are welcome